Joe's electric made 16 customer calls last week and 21 calls this week. how many calls must be made next in order to maintain anaverage of at least 19 calls per week for the three week period.
Answer by umnikmo(75)

Avr = 19 = (16+21+X)/3
from there :
3*19=16+21+X
X=20